Approximately how much bigger is a red blood cell than an E. coli cell
Answer #1

The size of Escherichia coli (E.coli) vary from 0.5 microns to 2 micrometer.

On an average it is m. ( 0.5 + 2/2 = 2.5/2 = 1.25 m )

The size of Red blood cell (RBC) vary from 6 microns to 9 micrometer.

On an average it is 7.5 m (6+9/2 = 15/2 = 7.5 m)

So, Size of RBC /Size of E.Coli = 7.5 m/1.25 m = 6

So, approximately RBC is 6 times larger than the Escherichia coli cell.